Jaleel invested in two mutual funds. One had a profit of 8%, the other 10%. He invested $1400 more in the 8% than the 10% fund. He earned $45 dollars more from the 8% fund than the 10% fund. How much did Jaleel invest in each fund?
Answer by stanbon(75887)

Jaleel invested in two mutual funds. One had a profit of 8%, the other 10%. He invested $1400 more in the 8% than the 10% fund. He earned $45 dollars more from the 8% fund than the 10% fund. How much did Jaleel invest in each fund?
------------------------------------
Equations:
E = T + 1400
0.08E = 0.10T + 45
------------------------
Substitute for "E" and solve for "T":
0.08(T + 1400) = 0.10T + 45
---
Multiply thru by 100:
8T + 8*1400 = 10T + 4500
2T = 6700
T = $3350 (amt. invested at 10%)
----
E = 3350+1400 = $4750 (amt. invested at 8%)
-----------------
cheers,
Stan H.